OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. @ryan123345

OpenStudy (ryan123345):

ok so the ''m'' is replaced with the 2 because u said 2 was the slope.

OpenStudy (ryan123345):

the letters that look like this are where u plug in the numbers

OpenStudy (ryan123345):

so in a coordinate u know that the first number is x and the second number is y right? (x,y)

OpenStudy (blurbendy):

use y - y1 = m(x - x1) y +2 = 2(x -3) y + 2 = 2x -6 y = 2x -8